# Autonomous Data Selection with Language Models for Mathematical Texts

### v2

@@ 4.2 Comparison with Baselines
Beyond complex reasoning, we extended our evaluation to assess how well the models adapted to other cognitive domains, such as commonsense reasoning, world knowledge, and reading comprehension. Table 4 encapsulates this multi-faceted performance evaluation. It’s noteworthy that while AutoDS did not universally top all categories, its overall average performance across diverse tasks (including all tasks shown in Table 3 and Table 4) shows the superiority of our method compared to other data selection methods. These outcomes strongly advocate for the AutoDS approach’s potential to advance language models in mathematical reasoning and beyond.
| Selection Method | MATH (5-shot) | GSM8K (5-shot) | BIG-Bench Hard (3-shot) |
| – (Mistral-7B Base) | 12.88 | 38.82 | 55.92 |
| Uniform (OpenWebMath) | 14.26 | 44.12 | 56.50 |
| DSIR | 12.30 | 42.00 | 55.97 |
| QuRating | 12.90 | 36.32 | 55.63 |
| AutoDS | 16.14 | 45.41 | 58.61 |
| H.S. | PIQA | W.G. | NQ | ARC-E | ARC-C | SciQ | LogiQA | BoolQ | |||
| Selection Method | (10) | (6) | (15) | (5) | (5) | (25) | (25) | (2) | (2) | (0) | Average |
| – (Mistral-7B Base) | 62.82 | 82.10 | 81.22 | 29.81 | 52.39 | 84.68 | 57.25 | 97.40 | 30.26 | 83.58 | 59.16 |
| Uniform (OpenWebMath) | 62.21 | 82.21 | 80.19 | 29.17 | 52.17 | 84.18 | 56.66 | 97.20 | 31.03 | 83.82 | 59.52 |
| DSIR | 63.10 | 81.94 | 81.37 | 29.22 | 52.62 | 84.72 | 57.25 | 97.30 | 30.26 | 73.76 | 58.59 |
| QuRating | 62.64 | 81.99 | 80.11 | 28.89 | 52.01 | 85.48 | 57.76 | 97.30 | 31.18 | 82.81 | 58.85 |
| AutoDS | 62.72 | 82.21 | 80.03 | 29.06 | 52.30 | 84.18 | 55.20 | 96.80 | 31.03 | 83.12 | 59.76 |
5 Related Work
Mathematical datasets and language models. The emergence of chain-of-thought prompting methodologies (Radford et al., 2019; Wei et al., 2022; Wang et al., 2022; Fu et al., 2022; Gao et al., 2023b; Yao et al., 2023; Zhang et al., 2023a; Gou et al., 2023) has been instrumental in harnessing and enhancing the reasoning capabilities inherent within language models. Our research, however, distinctly concentrates on the domain of continual pretraining with a focus on mathematical datasets. The creation of mathematical datasets has been critical in propelling AI models’ proficiency in mathematical comprehension and reasoning. Foundational contributions, such as the AMPS dataset by Hendrycks et al. (2021) and the Proof-Pile dataset by Azerbayev et al. (2023a), have provided capstones for models to systematically tackle mathematical problems and proofs. The Llemma model (Azerbayev et al., 2023b) builds upon this foundation, dedicating its efforts to the continual pretraining of language models with mathematical data especially the OpenWebMath dataset (Paster et al., 2023), aiming to refine their complex reasoning skills further. Nevertheless, the meticulous selection of mathematical data is still an area fraught with challenges.
Data selection in language modeling. The landscape of data selection in language modeling has seen a variety of approaches aimed at refining the quality and relevance of training data. Techniques have ranged from employing binary classifiers used by GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2023) to filter web data towards more formal sources like Wikipedia and books, to more nuanced strategies that consider the difficulty or domain-specificity of the data. For example, the Minerva model (Lewkowycz et al., 2022) used rule-based filtering for mathematical content, while DSIR (Xie et al., 2023b) applied importance resampling to align the data distribution with a target domain. Furthermore, DoReMi (Xie et al., 2023a) introduces a novel angle, optimizing domain weights with a proxy model to minimize worst-case excess loss across domains. However, the low inherent perplexity (entropy) in math-related and code-related corpora suggests that DoReMi might not be optimally suited for enhancing mathematical pretraining. Recently, Gunasekar et al. (2023); Li et al. (2023) demonstrated the utility of GPT-4 in annotating data quality for the Stack dataset (Kocetkov et al., 2022), subsequently using a random forest model for classification based on these annotations. Wettig et al. (2024) propose to train a reward model called Qurating for data selecting. Our work diverges from previous approaches by introducing a fully autonomous data selection method that leverages the intrinsic capabilities of language models without the need for human-generated (and AI-generated) annotations or external trained classifiers.
Data selection across various domains. The strategy of data selection transcends NLP tasks, extending its utility to a variety of domains, including vision and general domain adaptation. The Moore-Lewis technique, as introduced by Moore & Lewis (2010) and further refined by Axelrod (2017), exemplifies this approach by employing the cross-entropy differential between n-gram language models (LMs) tailored to specific targets and general corpora. Similarly, discrepancies in feature space and n-gram distributions have been effectively leveraged for data selection in domain adaptation scenarios, as evidenced by the work of Jiang & Zhai (2007), Liu et al. (2019), and Ruder & Plank (2017). Moreover, the significance of strategic data selection is equally acknowledged within the realm of computer vision, where methodologies aimed at optimizing training datasets have demonstrated substantial benefits. Notable contributions in this area include the pioneering curriculum learning framework by Bengio et al. (2009), the exploration of submodularity for efficient data selection by Wei et al. (2015), and recent advancements in prioritized data selection techniques by Coleman et al. (2019) and Mindermann et al. (2022).
@@ Appendix A Appendix for Examples
A.1 Web Subset
A.2 Code Subset
A.3 Arxiv Subset
@@ Appendix B More on Experiments
B.1 Prompts
B.2 Alternative Score functions

### v3

@@ 5 Experiments
Beyond complex reasoning, we extended our evaluation to assess how well the models adapted to other cognitive domains, such as commonsense reasoning, world knowledge, and reading comprehension. Table 3 encapsulates this multi-faceted performance evaluation. It’s noteworthy that while AutoDS did not universally top all categories, its average performance across diverse tasks (including all tasks shown in Table 2 and Table 3) shows the superiority of our method compared to other data selection methods. These outcomes strongly advocate for the AutoDS’s potential to advance language models in mathematical reasoning and beyond.
@@ Appendix A Appendix for Examples
Web Subset
Code Subset
Arxiv Subset
@@ Appendix B More on Experiments
More Experimental Results
| Model & Selection Method | HellaSwag (10-shot) | PIQA (6-shot) | WinoGrande (15-shot) |
| Gemma-2B Base | 48.30 | 70.67 | 60.54 |
| + Uniform (OpenWebMath) | 52.91 | 76.71 | 66.38 |
| + DSIR | 52.95 | 77.15 | 66.61 |
| + QuRating | 53.10 | 77.53 | 66.38 |
| + AutoDS | 52.82 | 77.42 | 66.61 |
| LLaMA2-7B Base | 58.88 | 79.43 | 75.85 |
| + Uniform (OpenWebMath) | 58.43 | 79.54 | 75.30 |
| + DSIR | 58.38 | 78.84 | 75.37 |
| + QuRating | 58.79 | 79.00 | 74.66 |
| + AutoDS | 58.28 | 78.18 | 74.51 |
| Mistral-7B Base | 62.82 | 82.10 | 81.22 |
| + Uniform (OpenWebMath) | 62.21 | 82.21 | 80.19 |
| + DSIR | 63.10 | 81.94 | 81.37 |
| + QuRating | 62.64 | 81.99 | 80.11 |
| + AutoDS | 62.72 | 82.21 | 80.03 |
Prompts
Alternative Score functions